A start-up publishing company estimates that the fixed costs of its first major project will be $190,000, the variable cost will be $18, and the selling price per book will be $34.

(a) How many books must be sold for this project to break even?

(b) Suppose the publishers wish to take a total of $40,000 in salary for this project. How many books must be sold to break even, and what is the break-even point, in dollars?
Answer #1

a) number of books to be sold for the projet to break even = fixed cost/contribution margin per unit = 190000/(34-18) = 11875

b) Now, number of books to be sold for the projet to break even = fixed cost/contribution margin per unit = (190000+40000)/(34-18) = 14375

Contribution margin ratio = Contributuion/sale price = ((34-18)/34)*100 = 47.05882353%

Break even point in sales dollars = Total fixed cost/Contribution margin ratio = (190000+40000)/47.05882353% = 488750 (Answer)